Using Exhibit 2, compute the percentage change in prices becwien (o) 1966 and 1969, (b) 1975 and 1986, and (c) 1990 and 1999. 8. 9, Assume the market basket contains 10X, 20Y, and 45Z. rices for goods X, Y, and Z are $1, $4, The current-year p and $6, respectively. $5, respectively. What is the CPI in the current year? The base-year prices are $1, $3. 10. If the CPI is 150 and nominal income is $100,000, what does real income equal?
media%2Fa73%2Fa73bb87d-2fc4-43e8-9644-e8
Answer #1

Question 8

(a)

CPI in 1966 = 32.4

CPI in 1969 = 36.7

Calculate the percentage change in price -

% change in price = [(CPI in 1969 - CPI in 1966)/CPI in 1966] * 100

% change in price = [(36.7 - 32.4)/32.4] * 100 = 13.27

The percentage change in price between 1966 and 1969 is 13.27%.

(b)

CPI in 1976 = 56.9

CPI in 1986 = 109.6

Calculate the percentage change in price -

% change in price = [(CPI in 1986 - CPI in 1976)/CPI in 1976] * 100

% change in price = [(109.6 - 56.9)/56.9] * 100 = 92.61

The percentage change in price between 1976 and 1986 is 92.61%.

(c)

CPI in 1990 = 130.7

CPI in 1999 = 166.6

Calculate the percentage change in price -

% change in price = [(CPI in 1999 - CPI in 1990)/CPI in 1990] * 100

% change in price = [(166.6 - 130.7)/130.7] * 100 = 27.47

The percentage change in price between 1990 and 1999 is 27.47%.

Question 10

CPI = 150

Nominal income = $100,000

Calculate the Real income -

Real income = Nominal Income * (100/CPI) = $100,000 * (100/150) = $66,666.67

The real income is $66,666.67